write a variation equation that has (3,11) as a solution then find y when x=12
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

y=kx
11=k*3
11/3=k
y=kx
y=11/3*12=44
y=(11/3)*x

Answer by stanbon(75887)

write a variation equation that has (3,11) as a solution then find y when x=12
-----------------------------
x = 3
y = 11
---------------
3x + 4y = 3*3 + 4*11
3x + 4y = 9+44
3x + 4y = 53
--------------------------------------------
The variation equation:
y = (-3/4)x + (53/4)
---
Find y when x = 12:
f(12) = (-3/4)*12 + (53/4)
f(12) = -9 + (53/4)
f(12) = 4.25
========================
Cheers,
Stan H.